# Analyzing the Performance Portability of SYCL across CPUs, GPUs, and Hybrid Systems with Protein Database Search

### v1

@@ Abstract
The high-performance computing (HPC) landscape is undergoing rapid transformation, with an increasing emphasis on energy-efficient and heterogeneous computing environments. This comprehensive study extends our previous research on SYCL’s performance portability by evaluating its effectiveness across a broader spectrum of computing architectures, including CPUs, GPUs, and hybrid CPU-GPU configurations from NVIDIA, Intel, and AMD. Our analysis covers single-GPU, multi-GPU, single-CPU, and CPU-GPU hybrid setups, using the SW# protein database search application as a case study. The results demonstrate SYCL’s versatility across different architectures, maintaining comparable performance to CUDA on NVIDIA GPUs while achieving similar architectural efficiency rates on most CPU configurations. Although SYCL showed excellent functional portability in hybrid CPU-GPU configurations, performance varied significantly based on specific hardware combinations. Some performance limitations were identified in multi-GPU and CPU-GPU configurations, primarily attributed to workload distribution strategies rather than SYCL-specific constraints. These findings position SYCL as a promising unified programming model for heterogeneous computing environments, particularly for bioinformatic applications.
@@ 1 Introduction
The high-performance computing (HPC) landscape is undergoing rapid transformation, with an increasing emphasis on energy-efficient and heterogeneous computing environments [navaux23]. GPUs are now the dominant accelerator in this field, offering incomparable performance for parallel workloads. Still, the roles of CPUs and the potential of hybrid CPU-GPU systems are also actively explored in computational research. On the one hand, GPU-based Systems on Chip (SoC) are becoming more popular every day. On the other hand, hybridization now extends to CPU architectures, with the emergence of heterogeneous designs like Intel’s Raptor Lake and Meteor Lake. Depending on the workload, these processors incorporate up to three different types of cores on the same die to optimize performance and energy efficiency. This architectural heterogeneity at the system and processor levels reflects the larger market dynamics and the ongoing evolution of computing architectures.
@@
This article evolves from our systematic exploration of SYCL’s capabilities in bioinformatics applications across heterogeneous computing platforms. In our initial work [Costanzo2022IWBBIO], we focused on migrating a CUDA-based, protein database search software (SW#db) to SYCL using Intel’s oneAPI ecosystem. This study demonstrated the tool’s effectiveness for code migration and it defined the basic cross-GPU-architecture portability of the migrated code, achieving comparable or slightly better efficiency rates compared to the native CUDA implementation. Building upon these findings, in [costanzo23] we expanded the study to evaluate SYCL’s performance portability across different GPU architectures, including single and multi-GPU configurations from different vendors. This research revealed that while CUDA and SYCL versions achieved similar performance on NVIDIA devices, SYCL reached higher architectural efficiency rates in most test cases from the other GPU vendors. Finally, in [supercompcost23] we completed the SYCL-migration process of SW# biological suite, analyzing its performance in a wide variety of biological test scenarios. Moreover, cross-SYCL-implementation portability was verified on several GPUs and some CPUs from a functional perspective.
This work extends our performance portability study to include (pure and hybrid) CPUs and CPU-GPU systems. The key contributions of this paper are:
@@
    An extension of our previous work on the functional and performance portability of SW# applications across different GPU architectures to now include Intel and AMD CPUs, as well as hybrid CPU-GPU configurations. On the one side, by investigating the performance characteristics of SYCL on these diverse platforms, this study aims to provide valuable insights into its potential as a unified programming model for the increasingly complex landscape of modern computing. On the other side, by expanding the study to include CPUs and hybrid systems, this research seeks to offer a more comprehensive understanding of how SYCL-based applications perform across the full spectrum of available computational resources.
@@
    An analysis of the trade-offs and synergies between CPU and GPU computations in hybrid configurations for a common bioinformatic workload.
@@ 2.1 CPUs, GPUs, and Programming Models for Heterogeneous Computing
The evolution of HPC has been significantly influenced by the emergence of general-purpose GPU programming. In 2007, NVIDIA introduced CUDA [CUDA_handson] alongside the Tesla GPU architecture, enabling non-graphics applications to harness GPU computational power. Although CUDA quickly became a prominent low-level programming model for GPU computing, its proprietary nature and exclusive compatibility with NVIDIA hardware led to the development of more versatile alternatives, such as OpenCL [OpenCL]. OpenCL offers a level of abstraction comparable to that of CUDA while supporting multiple devices and vendors.
@@
2.2 Smith-Waterman Protein Database Search
A protein database search is a bioinformatics process where researchers or scientists look for specific protein sequences or related information within a large collection of known protein data. The search begins with a query, which is usually an unknown protein sequence, which is compared against each sequence from a known protein database. The comparison method used for each pair of sequences is the SW algorithm, which provides the optimal local alignment between two sequences [Smith1981]. This method is based on a dynamic programming approach and is highly sensitive since it explores all possible alignments between the sequences.
@@ 2.2 Smith-Waterman Protein Database Search
The exploitation of parallelism is limited by the data dependencies present in the calculation of $H_{i,j}$. Because of each alignment is independent, the most commonly used approach in this case is inter-task parallelism. This scheme is based on computing multiple matrices simultaneously and is illustrated in Fig. 1.
@@
The SW algorithm runs in quadratic time and space to compute optimal alignment. However, it can be calculated in linear space complexity because computing optimal alignment scores does not require storing the full similarity matrix. Similarity database search takes advantage of this feature since optimal alignment only makes sense for very similar sequences. Therefore, all alignment scores are calculated first and optimal alignments are computed only for top-ranked database sequences.
@@ 3.1 Case-Study Applications
Two GPU-accelerated implementations of protein database search were considered for the performance portability evaluation:
@@
    CUDA: The SW# framework represents a CUDA-implemented solution for biological sequence alignment that prioritizes memory efficiency. Operating as both a standalone program and an integrated library, it enables protein and DNA sequence comparisons through pairwise alignment and database searches. The framework provides customizable alignment parameters, including Smith-Waterman algorithm options, gap penalty specifications, and scoring matrix selection. Through CPU-GPU coordination, SW# achieves good performance by implementing a dynamic load balancing system that assigns tasks based on sequence length characteristics. The parallelization architecture primarily leverages GPU capabilities through both inter-task and intra-task approaches. GPU processing follows a dual-kernel strategy: shorter database sequences are processed via inter-task parallelism in a “short kernel”, while longer sequences utilize intra-task parallelism in a “long kernel.” For multi-GPU configurations, SW# employs an adaptive execution protocol. When GPU count exceeds query sequence numbers, all devices synchronously process different database segments for a single query. Alternatively, with more queries than GPUs, each GPU independently processes distinct queries against the complete database [swsharp, swsharpdb].
@@ 3.2 Performance Model
where Clock_rate denotes the device’s clock frequency,
Throughput represents the number of instructions the device can execute per clock cycle, and Lanes indicates the quantity of SIMD vector lanes. Subsequently, it is imperative to enumerate the instructions issued during each cell update of the similarity matrix. In the sequence alignment context, the predominant performance metric is Cell Updates Per Second (CUPS) [Costanzo2022IWBBIO]. Consequently, the theoretical peak performance of any device can be modeled using Equation 5:
@@ 3.2.2 Architectural features on NVIDIA GPUs
The # Cores in an NVIDIA GPU refers to the number of Streaming Multiprocessors. CUDA does not strictly follow a SIMD execution model but it adopts a similar one denoted as the SIMT model. A warp is composed of a group of 32 threads that execute the same instruction stream. According to [lan2017swhybrid], “a warp in SIMT is equivalent to a vector in SIMD, and a thread in SIMT is equivalent to a vector lane in SIMD”. The instruction throughput depends on the CUDA Compute Capability (CC) of each NVIDIA GPU  [^7] https://docs.nvidia.com/cuda/cuda-c-programming-guide/#maximize-instruction-throughput.
@@ 4.1 Experimental Design
The experiments were carried out on 12 GPUs, including 6 NVIDIA dGPUs, 1 AMD dGPU, 1 AMD iGPU, 3 Intel iGPUs, and 1 Intel dGPU. In addition, 8 Intel CPUs (from different segments) and 1 AMD CPU were included in the study. The specific details of these platforms can be found in Table 1. The oneAPI and CUDA versions used were 2022.1.0 and 11.7, respectively. For both CUDA and SYCL, the optimization flag -O3 was used during compilation. To run SYCL code on NVIDIA and AMD GPUs, several modifications had to be made to the build process, as SYCL is not supported by default on these platforms [^9] https://intel.github.io/llvm-docs/GetStartedGuide.html but Codeplay recently announced free binary plugins [^10] https://codeplay.com/portal/blogs/2022/12/16/bringing-nvidia-and-amd-support-to-oneapi.html to support it. After these modifications, it was possible to run DPC++ code on an NVIDIA GPU using the Clang++ compiler (16.0).
@@
Performance evaluation was carried out by searching for 20 query protein sequences against the well-known Environmental Non-Redundant database (Env. NR) (2021_04 Release), which contains 995210546 amino acid residues in 4789355 sequences, with a maximum length of 16925. Query sequences were selected from the Swiss-Prot database  [^11] Swiss-Prot:  https://www.uniprot.org/downloads, with lengths ranging from 144 to 5478. The access numbers for these queries are: P02232, P05013, P14942, P07327, P01008, P03435, P42357, P21177, Q38941, P27895, P07756, P04775, P19096, P28167, P0C6B8, P20930, P08519, Q7TMA5, P33450, and Q9UKN1. Besides, SW# was configured with BLOSUM62 as the substitution matrix, and 10/2 as the penalty for insertion/extension. Last, each test was executed 20 times and the performance was determined based on the average of these multiple runs (to minimize deviations).
@@ 4.2 Single-GPU Performance and Portability Results
A comprehensive analysis of SYCL and CUDA implementations was conducted on various NVIDIA, AMD, and Intel GPUs to evaluate the performance and portability of SYCL across different GPU architectures. Table LABEL:tab:arch-eff provides a detailed comparison of the performance and the architectural efficiency of both implementations on each platform.
On NVIDIA GPUs, CUDA and SYCL demonstrated comparable performance and efficiency values. As anticipated, more powerful GPUs achieved higher GCUPS values. The architectural efficiency values ranged from 37% to 52%. In particular, while the RTX 3090 GPU exhibited the highest GCUPS value, the RTX 2070 GPU proved to be the most efficient considering architectural usage.
For AMD and Intel GPUs, only SYCL results are presented, as CUDA exclusively supports NVIDIA hardware. This limitation underscores the superior portability of SYCL compared to CUDA. The performance of the SYCL implementation on AMD and Intel GPUs presents a mixed picture. While AMD’s RX 6700 XT dGPU achieved efficiency rates comparable to NVIDIA GPUs (51.7%), the AMD Vega 6 iGPU showed considerably lower performance with only 21.3% efficiency rate. For Intel GPUs, a notable dichotomy emerged between their graphic architectures. Intel’s Gen-based iGPUs (UHD 630 and UHD 770) demonstrated impressive architectural efficiency, reaching up to 75.7%, significantly surpassing the baseline established by NVIDIA GPUs. However, Intel’s Xe-based GPUs (Arc A770 dGPU and Xe-LPG 128EU iGPU) exhibited markedly lower performance, with architectural efficiencies down to 18%, respectively. This suboptimal performance can be attributed to ongoing driver optimization challenges in Intel’s Arc generation, a factor that has contributed to Intel’s dGPU market share dropping below 1% in 2024 [norem24]. These performance limitations with Xe-based GPUs will likely impact the overall performance portability metrics for Intel GPUs in this study. Further investigation, including detailed code profiling, is planned to understand these performance discrepancies and their relationship to hardware architecture and driver maturity.

### v3

# Analyzing the Performance Portability of SYCL across CPUs, GPUs, and Hybrid Systems with SW Sequence Alignment

@@ Abstract
The high-performance computing (HPC) landscape is undergoing rapid transformation, with an increasing emphasis on energy-efficient and heterogeneous computing environments. This comprehensive study extends our previous research on SYCL’s performance portability by evaluating its effectiveness across a broader spectrum of computing architectures, including CPUs, GPUs, and hybrid CPU-GPU configurations from NVIDIA, Intel, and AMD. Our analysis covers single-GPU, multi-GPU, single-CPU, and CPU-GPU hybrid setups, using two common, bioinformatic applications as a case study. The results demonstrate SYCL’s versatility across different architectures, maintaining comparable performance to CUDA on NVIDIA GPUs while achieving similar architectural efficiency rates on AMD and Intel GPUs in the majority of cases tested. SYCL also demonstrated remarkable versatility and effectiveness across CPUs from various manufacturers, including the latest hybrid architectures from Intel. Although SYCL showed excellent functional portability in hybrid CPU-GPU configurations, performance varied significantly based on specific hardware combinations. Some performance limitations were identified in multi-GPU and CPU-GPU configurations, primarily attributed to workload distribution strategies rather than SYCL-specific constraints. These findings position SYCL as a promising unified programming model for heterogeneous computing environments, particularly for bioinformatic applications.
@@
[…]
[L] © $<$2025$>$. This manuscript version is made available under the CC-BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/
The published journal article is available at: https://doi.org/10.1016/j.future.2025.107838
@@ 1 Introduction
The parallel processing landscape is undergoing rapid transformation, with an increasing emphasis on energy-efficient and heterogeneous computing environments [navaux23]. GPUs are now the dominant accelerator in this field, offering incomparable performance for parallel workloads. Still, the roles of CPUs and the potential of hybrid CPU-GPU systems are also actively explored in computational research. On the one hand, GPU-based Systems on Chip (SoC) are becoming more popular every day. On the other hand, hybridization now extends to CPU architectures, with the emergence of heterogeneous designs like Intel’s Raptor Lake and Meteor Lake. Depending on the workload, these processors incorporate up to three different types of cores on the same die to optimize performance and energy efficiency. This architectural heterogeneity at the system and processor levels reflects the larger market dynamics and the ongoing evolution of computing architectures.
@@
This article evolves from our systematic exploration of SYCL’s capabilities in bioinformatics applications across heterogeneous computing platforms. In our initial work [Costanzo2022IWBBIO], we focused on migrating a CUDA-based, protein database search software (SW#db) to SYCL using Intel’s oneAPI ecosystem. This study demonstrated the tool’s effectiveness for code migration and it defined the basic cross-GPU-architecture portability of the migrated code, achieving comparable or slightly better efficiency rates compared to the native CUDA implementation. Building upon these findings, in [costanzo23] we expanded the study to evaluate SYCL’s performance portability across different GPU architectures, including single and multi-GPU configurations from different vendors. This research revealed that while CUDA and SYCL versions achieved similar performance on NVIDIA devices, SYCL reached higher architectural efficiency rates in most test cases from the other GPU vendors. Finally, in [supercompcost23] we completed the SYCL-migration process of SW# biological suite to support pairwise alignment and different alignment algorithms. Moreover, its performance in a wide variety of biological test scenarios was analyzed and cross-SYCL-implementation portability was verified on several GPUs and some CPUs from a functional perspective.
This work extends our performance portability study to include pairwise sequence alignment and (pure and hybrid) CPUs and CPU-GPU systems. The key contributions of this paper are:
@@
    An extension of our previous work on the functional and performance portability of SW# applications across different GPU architectures to now include pairwise sequence alignment and Intel and AMD CPUs, as well as hybrid CPU-GPU configurations. On the one side, by investigating the performance characteristics of SYCL on these diverse platforms, this study aims to provide valuable insights into its potential as a unified programming model for the increasingly complex landscape of modern computing. On the other side, by expanding the study to include CPUs and hybrid systems, this research seeks to offer a more comprehensive understanding of how SYCL-based applications perform across the full spectrum of available computational resources.
@@
    An analysis of the trade-offs and synergies between CPU and GPU computations in hybrid configurations for two common bioinformatic workloads.
@@ 2.1 CPUs, GPUs, and Programming Models for Heterogeneous Computing
The evolution of high-performance computing (HPC) has been significantly influenced by the emergence of general-purpose GPU programming. In 2007, NVIDIA introduced CUDA [CUDA_handson] alongside the Tesla GPU architecture, enabling non-graphics applications to harness GPU computational power. Although CUDA quickly became a prominent low-level programming model for GPU computing, its proprietary nature and exclusive compatibility with NVIDIA hardware led to the development of more versatile alternatives, such as OpenCL [OpenCL]. OpenCL offers a level of abstraction comparable to that of CUDA while supporting multiple devices and vendors.
@@
2.2 Smith-Waterman Sequence Alignment
A key process in bioinformatics and computational biology is sequence alignment, which aims to identify regions of similarity between sequences to uncover their structural, functional, and evolutionary relationships [PSAreview]. The Smith-Waterman (SW) algorithm is commonly employed to determine the optimal local alignment between two sequences [Smith1981]. This technique relies on dynamic programming and is highly sensitive, as it examines all possible alignments between the sequences.
@@ 2.2 Smith-Waterman Sequence Alignment
SW is usually employed to compute: (a) pairwise alignments of two very long DNA sequences, and (b) database similarity search (one-to-many), involving shorter protein sequences. As SW runs in quadratic time to compute optimal alignment, widely adopted parallelization approaches are intra-sequence (a single matrix is calculated and all Processing Elements (PEs) work collaboratively) and inter-sequence (simultaneously calculating multiple matrices without communication between the PEs). Fig. 1 illustrates both approaches.
@@ 3.1 Case-Study Applications
Two GPU-accelerated implementations of SW sequence alignment were considered for the performance portability evaluation:
@@
    CUDA: The SW# framework represents a CUDA-implemented solution for biological sequence alignment that prioritizes memory efficiency. Operating as both a standalone program and an integrated library, it enables protein and DNA sequence comparisons through pairwise alignment and database searches. The framework provides customizable alignment parameters, including Smith-Waterman algorithm options, gap penalty specifications, and scoring matrix selection. Through CPU-GPU coordination, SW# achieves good performance by implementing a dynamic load balancing system that assigns tasks based on sequence length characteristics. The parallelization architecture primarily leverages GPU capabilities through both inter-task and intra-task approaches. GPU processing follows a dual-kernel strategy: shorter database sequences are processed via inter-task parallelism in a short kernel, while longer sequences utilize intra-task parallelism in a long kernel. For multi-GPU configurations, SW# employs an adaptive execution protocol. When the GPU count exceeds query sequence numbers, all devices synchronously process different database segments for a single query. Alternatively, with more queries than GPUs, each GPU independently processes distinct queries against the complete database [swsharp, swsharpdb].
@@ 3.2 Performance Model
where Clock_rate denotes the device’s clock frequency, Throughput represents the number of instructions the device can execute per clock cycle, and Lanes indicates the quantity of SIMD vector lanes. Subsequently, it is imperative to enumerate the instructions issued during each cell update of the similarity matrix. In the sequence alignment context, the predominant performance metric is Cell Updates Per Second (CUPS), representing the time for a complete computation of one cell in the similarity matrix (H). The GCUPS (billion CUPS) value is calculated by:
$GCUPS=\frac{Q\times D}{t\times 10^{9}}$ (5)
where $|Q|$ is the total number of residues in the query sequence, $|D|$ is the total number of residues in the database and t is the runtime in seconds [RucciBDAG2016]. Consequently, the theoretical peak GCUPS of any device can be modeled using Equation 6:
@@ 3.2.2 Architectural features on NVIDIA GPUs
The # Cores in an NVIDIA GPU refers to the number of Streaming Multiprocessors. CUDA does not strictly follow a SIMD execution model but it adopts a similar one denoted as the SIMT model. A warp comprises a group of 32 threads that execute the same instruction stream. According to [lan2017swhybrid], “a warp in SIMT is equivalent to a vector in SIMD, and a thread in SIMT is equivalent to a vector lane in SIMD”. The instruction throughput depends on the CUDA Compute Capability (CC) of each NVIDIA GPU  [^7] https://docs.nvidia.com/cuda/cuda-c-programming-guide/#maximize-instruction-throughput.
@@ 4.1 Experimental Design
The experiments were carried out on 12 GPUs, including 6 NVIDIA dGPUs, 1 AMD dGPU, 1 AMD iGPU, 3 Intel iGPUs, and 1 Intel dGPU. In addition, 8 Intel CPUs (from different segments) and 1 AMD CPU were included in the study. The specific details of these platforms can be found in Table 1. The oneAPI and CUDA versions used were 2022.1.0 and 11.7, respectively. For both CUDA and SYCL, the optimization flag -O3 was used during compilation. To run SYCL code on NVIDIA and AMD GPUs, several modifications had to be made to the build process, as SYCL is not supported by default on these platforms [^9] https://intel.github.io/llvm-docs/GetStartedGuide.html but Codeplay recently announced free binary plugins [^10] https://codeplay.com/portal/blogs/2022/12/16/bringing-nvidia-and-amd-support-to-oneapi.html to support it. After these modifications, it was possible to run DPC++ code on both NVIDIA and AMD GPUs using the Clang++ compiler (16.0).
@@
Performance evaluation was carried out using real biological data for the two application benchmarks:
  • 1.
    Protein database search: searching 20 query protein sequences against the well-known Environmental Non-Redundant database (Env. NR) (2021_04 Release), which contains 995210546 amino acid residues in 4789355 sequences, with a maximum length of 16925. Query sequences were selected from the Swiss-Prot database  [^11] Swiss-Prot:  https://www.uniprot.org/downloads, with lengths ranging from 144 to 5478. The access numbers for these queries are: P02232, P05013, P14942, P07327, P01008, P03435, P42357, P21177, Q38941, P27895, P07756, P04775, P19096, P28167, P0C6B8, P20930, P08519, Q7TMA5, P33450, and Q9UKN1. Besides, SW# was configured with BLOSUM62 as the substitution matrix, and 10/2 as the penalty for insertion/extension.
  • 2.
    Pairwise alignment: computing five different DNA alignments. Table 3 presents the accession numbers and sizes of the DNA sequences used. The score parameters were configured as +1 for match, -3 for mismatch, -5 for gap open, and -2 for gap extension.
Query sequence Database sequence
Accession Size Accession Size
Matrix Size
NC_000898 162K NC_007605 172K 28M
NC_003064.2 543K NC_000914.1 536K 291M
CP000051.1 1M AE002160.2 1M 1G
BA000035.2 3M BX927147.1 3M 9G
NC_005027.1 7M NC_003997.3 5M 35G
Table 3: DNA sequence information used in the tests
Last, each test was executed 20 times and the performance was determined based on the average of these multiple runs (to minimize deviations).
@@ 4.2 Single-GPU Performance and Portability Results
A comprehensive analysis of SYCL and CUDA implementations was conducted on various NVIDIA, AMD, and Intel GPUs to evaluate the performance and portability of SYCL across different GPU architectures.
4.2.1 Protein Database Search
Table LABEL:tab:arch-eff provides a detailed comparison of the performance and the architectural efficiency of CUDA and SYCL implementations on each platform for protein database search. On NVIDIA GPUs, CUDA and SYCL demonstrated comparable performance and efficiency values. As anticipated, more powerful GPUs achieved higher GCUPS values. The architectural efficiency values ranged from 37% to 52%. In particular, while the RTX 3090 GPU exhibited the highest GCUPS value, the RTX 2070 GPU proved to be the most efficient considering architectural usage.
For AMD and Intel GPUs, only SYCL results are presented, as CUDA exclusively supports NVIDIA hardware. This limitation underscores the superior portability of SYCL compared to CUDA. The performance of the SYCL implementation on AMD and Intel GPUs presents a mixed picture. While AMD’s RX 6700 XT dGPU achieved efficiency rates comparable to NVIDIA GPUs (51.7%), the AMD Vega 6 iGPU showed considerably lower performance with only 21.3% efficiency rate. For Intel GPUs, a notable dichotomy emerged between their graphic architectures. Intel’s Gen-based iGPUs (UHD 630 and UHD 770) demonstrated impressive architectural efficiency, reaching up to 75.7%, significantly surpassing the baseline established by NVIDIA GPUs. However, Intel’s Xe-based GPUs (Arc A770 dGPU and Xe-LPG 128EU iGPU) exhibited markedly lower performance, with architectural efficiencies down to 23.4%, respectively. This suboptimal performance can be attributed to ongoing driver optimization challenges in Intel’s Arc generation, a factor that has contributed to Intel’s dGPU market share dropping below 1% in 2024 [norem24]. These performance limitations with Xe-based GPUs will likely impact the overall performance portability metrics for Intel GPUs in this study. Further investigation, including detailed code profiling, is planned to understand these performance discrepancies and their relationship to hardware architecture and driver maturity.